Cobalt-60, with a half-life of 5.3 years, has a number of

applications, including medical therapy and the sterilization

of medical tools. Determine the mass of a 50 g sample that

would remain after

(a) 6 months

(b) 5 years


1
Expert's answer
2022-03-15T20:13:45-0400

"N=N_0(\frac{1}{2})^{\frac{t}{t_{1/2}}}"


a)

"N=50\ g\times(\frac{1}{2})^{\frac{0.5\ years}{5.3\ years}}=47\ g"


Answer: 47 g


b)

"N=50\ g\times(\frac{1}{2})^{\frac{5\ years}{5.3\ years}}=26\ g"


Answer: 26 g